OpenStudy (deadpooi):

What is the value of x? 3 • (–4) • x = 0 x =

OpenStudy (misssmartiez):

3 x -4 x - = 0. 3 x -4 = -12. -12 x - = 0?

OpenStudy (misssmartiez):

Hint: 0 x v will always result to 0.

rebeccaxhawaii (rebeccaxhawaii):

as ms smarts says -12x=0 to get x by itself we would divide -12 to both sides

OpenStudy (misssmartiez):

Thanks, reb.

OpenStudy (deadpooi):

-6? @MissSmartiez and @rebeccaxhawaii

rebeccaxhawaii (rebeccaxhawaii):

nope what is 0/-12

OpenStudy (deadpooi):

0?

OpenStudy (misssmartiez):

As I said.

OpenStudy (misssmartiez):

correct

OpenStudy (deadpooi):

Oh so were diving it by both sides.!!!

OpenStudy (deadpooi):

Thank you all